Primitives of Sequential and Parallel

Computation

M. Forsell, V. Leppänen, M. Penttonen

Report A/1998/6

ISBN 951-781-608-1 ISSN 0787-6416

UNIVERSITY OF KUOPIO Department of Computer Science

and Applied Mathematics

P.O.Box 1627, FIN-70211 Kuopio, FINLAND